OpenStudy (anonymous):

What IS the percent of 16 out of 30 what is 16 0ut 0f 30 in a decimal What Is 16 0ut Of 30 in A Angle The Degrees Of The Angle is 360 HELP ME PLEASE !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

percentage: 16/30*100=53.33% decimal: 16/30 =.5333 16/30 of 360=16/30*360=16*12=192 degrees